OpenStudy (anonymous):

the product of 2 negative integers is 36. the second integer is 5 more than the first. find the integer

OpenStudy (anonymous):

Let x = first integer x + 5 = second integer x(x+5) = 36 Now you solve that. Use Distributive Law, set equal to zero, factor, and solve.

OpenStudy (anonymous):

so x^2 is 36?

OpenStudy (anonymous):

NO.

OpenStudy (anonymous):

I said, x(x+5) = 36 Can you begin solving that? What should be your FIRST step?

OpenStudy (anonymous):

im sorry im just confused and i have no idea what so ever how to do this

OpenStudy (anonymous):

If you have no idea as to how to start solving that equation, then relearn the stepsin solving equations.